BTGE2033 ENGINEERING STATISTICS Q4. (Continued) (d) A stochastic process can be in one of three possible states: state 1, state 2 and state 3. The transition probability from state i to state j is represented by the entry at row i column j of the following 1-step transition matrix 0.4 0.3 0.3 P0 0.4 0.6 0.6 0 0.4, (0.7, 0, 0.3) and the probability The initial probability distribution of the states is vo distribution of the states after n steps of transition is represented by Vn (3 marks) (i) Calculate v2 (ii) Calculate the steady-state probability distribution of the states, ve, which is (4 marks) attained after a very big number of transitions. Total: 25 marks
Answer #1

(i)

0.4 0.3 0.3 P= 0 0.4 0.6 0.6 0 0.4ノ

0.4 0.3 0.3/0.4 0.3 0.3/0.34 0.24 0.42 P2= 0 0.4 0.6 0 0.4 0.6 = 0.36 0.16 0.48 0.6 0 0.4(0.6 0 0.40.48 0.18 0.34/

0.34 0.24 0.42 v2P2 -(0.7 0 0.3) 0.360.16 0.48(0.382 0.222 0.396) 0.48 0.18 0.34/

(ii)

For steady state probability distributions,

ひ

Let v_{infty } = (a, b, c)

then,

a + b + c = 1

0.4a + 0.6c = a => 0.6c = 0.6a => a = c

0.3a + 0.4b = b => 0.3a = 0.6b => a = 2b => c = 2b

0.3a + 0.6b + 0.4c = c => 0.3a + 0.6b - 0.6c = 0

Now,

a + b + c = 1

=> 2b + b + 2b = 1

=> 5b = 1

=> b = 0.2

a = c = 2b = 2 * 0.2 = 0.4

Thus, the steady state probability distributions is,

v_{infty } = (0.4, 0.2, 0.4)